# RLSA-PFL: Robust Lightweight Secure Aggregation with Model Inconsistency Detection in Privacy-Preserving Federated Learning

### v2

@@ 5.2 Main Scheme
The intermediate servers and aggregator collect masked models from at least $t$ users, where $t$ is the threshold required for each training round. This threshold $t$ is critical for maintaining the security of the masked models. For example, if only two users participate in a round, one user’s local model could be exposed to the other, increasing the risk of collusion. Therefore, the larger the value of $t$, the more secure the system becomes. The intermediate servers and the aggregator wait for a certain timeout period to gather enough masked models before aborting the process if an insufficient number is collected. Users are free to drop out at any time.
@@
Afterward, the aggregated global model $\theta$ is shared with all active participants for the next round of training. Simultaneously, the aggregator sends the tuple $V$ to each intermediate server. The intermediate server forwards a part of tuple $V$ to the users in $\mathrm{I}$ for verification. If the verification of the tuples is successful, the users participate in the next round; otherwise, they abort.
Figure 3 presents both the semi-honest and malicious model versions of the protocol. In the semi-honest model, where all entities follow the protocol honestly, the use of signatures and a PKI can be avoided.
6 Security Analysis
In this section, we present the security claims and their respective proofs for our proposed scheme discussed in Section 5.2.
6.1 Semi-Honest Model
In this section, we demonstrate that even if a threshold number of users, denoted as $t_{c}$, and intermediate servers, denoted as $t_{f}$, collude among themselves or with the aggregator, they cannot learn about the remaining genuine users’ local weighted models. We follow a similar security model to those in [12] and [17].
We consider three scenarios: first, a subset of users are dishonest and collude with adversaries or among themselves, while the intermediate servers and the aggregator remain honest. Second, a subset of intermediate servers are dishonest and can collude, while the users and the aggregator are honest. Third, a subset of users and intermediate servers, along with the aggregator, are dishonest and can collude. Our goal is to show that if fewer than $t_{c}$ (where $t_{c}>2$) and $t_{f}$ (where $t_{f}>1$) entities are compromised, our scheme still protects the individual locally trained models of the remaining genuine users. We assume that $\mathbb{U}$ and $\mathbb{F}$ be the total number of users and intermediate servers, respectively. We also assume that $\mathcal{U}_{i}$ be the number of participating users in the $i^{th}$ round, and $x_{\mathbb{U}}$ is the locally trained models of the users $\mathbb{U}$.
@@ 6.1 Semi-Honest Model
Proof 2
In our scheme, users randomly choose their input values (i.e., random secret keys) to mask the local model parameters, similar to a one-time pad. Each user’s input values are independent of those of other users. Since the aggregator and intermediate servers are considered honest entities, the combined view of the users in $C_{c}$ will not reveal any useful information about the input values of users not in $C_{c}$. Additionally, the honest-but-curious users in $C_{c}$ only receive the identities of the honest users not in $C_{c}$. This enables the simulator $\mathsf{Sim_{C}^{\mathbb{U},t_{c},\lambda}}$ to use dummy input values for the honest users not in $C_{c}$ while keeping the combined views of the users in $C_{c}$ identical to that of $\mathsf{Real_{C}^{\mathbb{U},t_{c},\lambda}}$.
Theorem 3
@@
Proof 4
Theorem 5
@@
Proof 6
We present the detailed security proof in the Appendix A.
6.2 Malicious Model
@@
In this section, we follow the standard security model as in [12] and [17] to demonstrate that our scheme is secure against active malicious attackers.
@@
In this security model, we need to consider three cases. The first case is the Sybil Attack, where the adversaries can simulate a specific honest user $u_{i}$ to get its inputs. Please note that our scheme is resistant to this attack, as it uses digital signatures to verify the origin of the messages. The second case involves the aggregator sending different lists of participating users in a round to the honest intermediate servers. This could lead to the disclosure of the local weighted models of the honest users. Our scheme detects this type of attack during the model parameter verification phase, where each user can verify the list of participating users in a round received from the fog servers. Any difference in the lists will indicate an attack.
@@
The third case involves an aggregator dishonestly dropping honest users at any round. We use the Random Oracle model to prove that our scheme is secure against any malicious dropout of honest users by adversaries. Let’s consider $M_{c}$ as a probabilistic polynomial-time algorithm representing the next message function of participants in $C$, which enables users in $C$ to dynamically select their inputs at any round of the protocol and the list of participating users. The Random Oracle can output the sum of a dynamically selected subset of honest clients for the simulator $\mathrm{Sim}$, which is indistinguishable from the combined view of adversaries in the real protocol execution $\mathsf{Real}(M_{c})$. There are three possible scenarios for this type of active attack: first, the malicious users can collude among themselves; second, the malicious intermediate servers can collude among themselves; and finally, the malicious users and intermediate servers can collude with the aggregator. We present the following three theorems to demonstrate our security proofs.
TABLE II: Running Time in Microseconds for User, Intermediate Server, and Aggregator in Different Phases Under Semi-Honest Model with a Vector Size of $50$k, $64$-bit Length, and $10$ Intermediate Servers per Round.
User Drop Outs $0\%$ $10\%$ $20\%$ $30\%$
No. of Users $500$ $1000$ $500$ $1000$ $500$ $1000$ $500$ $1000$
Masking Time at User $4031$ $4025$ $3915$ $4096$ $4099$ $4085$ $3921$ $4043$
Partial Aggregation Time at Intermediate Server $26014.9$ $51575.8$ $23951.7$ $46535.6$ $21353$ $43355.2$ $18348.5$ $35440.2$
Final Aggregation Time at Aggregator $26816$ $52422$ $24755$ $47422$ $22109$ $44262$ $19101$ $36284$
Initiation of Global Model Parameter Verification Time at Aggregator $3924$ $3936$ $3938$ $3931$ $3944$ $3930$ $3881$ $3918$
Global Model Parameter Verification Time at User $3748$ $3746$ $3754$ $3764$ $3757$ $3739$ $3704$ $3749$
Aggregator Total Running Time $30740$ $56358$ $28693$ $51353$ $26053$ $48192$ $22982$ $40202$
Intermediate Server Total Running Time $26343$ $52218$ $24254$ $47135$ $21629$ $43908$ $18590$ $35926$
User Total Running Time $7779$ $7771$ $7669$ $7860$ $7853$ $7824$ $7625$ $7792$
@@
TABLE III: Running Time in Microseconds for User, Intermediate Server, and Aggregator in Different Phases under the Malicious Model with a Vector Size of $50k$, 64-bit Length, and 10 Intermediate Servers per Round.
User Drop Outs $0\%$ $10\%$ $20\%$ $30\%$
No. of Users $500$ $1000$ $500$ $1000$ $500$ $1000$ $500$ $1000$
Masking Time at User $7671$ $7702$ $7654$ $7739$ $7681$ $7721$ $7679$ $7840$
Partial Aggregation Time at Intermediate Server $29025$ $55369.2$ $25993$ $52227$ $23671$ $43868$ $21049$ $38870$
Final Aggregation Time at Aggregator $55445$ $82123$ $52090$ $78668$ $50075$ $70064$ $47305$ $65132$
Initiation of Global Model Parameter Verification Time at the Aggregator $7671$ $7702$ $7654$ $7739$ $7681$ $7721$ $7679$ $7840$
Global Model Parameter Verification Time at User $3777$ $3797$ $3774$ $3810$ $3782$ $3807$ $3783$ $3817$
Aggregator Total Running Time $63116$ $89825$ $59744$ $86407$ $57756$ $77785$ $54984$ $72972$
Intermediate Server Total Running Time $1286470$ $2534286$ $1145840$ $2313726$ $1024000$ $2044189$ $892556$ $1786718$
User Total Running Time $49112$ $49368$ $49014$ $49678$ $49220$ $49515$ $49001$ $49498$
(a) Intermediate Server Average Running Time in Microseconds per Round
(b) Aggregator Running Time in Microseconds per Round
@@
Figure 4: Average Running Time of an Intermediate Server and the Aggregator with No User Dropout while Varying the Number of Users in the Semi-Honest Model per Round
(a) Intermediate Server Average Running Time in Seconds per Round
(b) Aggregator Running Time in Microseconds per Round
@@
Figure 5: Average Running Time of an Intermediate Server and the Aggregator with No User Dropout while Varying the Number of Users in a Malicious Model per Round
(a) An Intermediate Server Average Running Time in Microseconds per Round
(b) Aggregator Running Time in Microseconds per Round
@@
Figure 6: Average Running Time of an Intermediate Server and the Aggregator with No User Dropout while Varying the Size of Vectors in a Semi-Honest Model per Round
(a) An Intermediate Server Average Running Time in Microseconds per Round
(b) Aggregator Running Time in Microseconds per Round
@@
Figure 7: Average Running Time of an Intermediate Server and the Aggregator with No User Dropout while Varying the Size of Vectors in a Malicious Model per Round
Theorem 7
@@
Proof 8
This proof is similar to that of Theorem 1, as the users in $C_{c}$ gain no knowledge about $x_{\mathbb{U}\setminus C_{c}}$. Consequently, the simulator $\mathsf{Sim}$ can assign real inputs to the dishonest users while assigning dummy inputs to the remaining users, accurately replicating the perspectives of the users in $C_{c}$. Thus, the joint views of the users in $C_{c}$ in the simulation are indistinguishable from those in $\mathsf{Real}$.
Theorem 9
@@
Proof 10
This proof is similar to that of Theorem 3, as the intermediate servers in $C_{f}$ gain no knowledge about $x_{\mathbb{F}\setminus C_{f}}$ apart from the list of participants. Consequently, the simulator $\mathsf{Sim}$ can assign real inputs to the dishonest intermediate servers while assigning dummy inputs to the remaining intermediate servers, accurately replicating the perspectives of the intermediate servers in $C_{f}$. Thus, the joint views of the intermediate servers in $C_{f}$ in the simulation are indistinguishable from those in $\mathsf{Real}$.
Theorem 11
@@
Proof 12
We present the detailed security proof in the Appendix B.
7 Performance Evaluation
In this section, we first present a theoretical performance evaluation, covering computation and communication complexity, as well as dropout resilience, followed by the experimental results of our scheme.
7.1 Theoretical Analysis
We structure our theoretical analysis into two main categories: computation and communication complexity, and security and functionality. We compare our scheme with closely related works, including SecAgg [12], SecAgg+ [13], TurboAgg [30], FastSecAgg [21], and EDRAgg [17].
Section 7.1.1 provides a detailed analysis of the computation and communication costs for users, intermediate servers, and the main aggregator, highlighting the worst-case scenarios. While our approach introduces an intermediate server layer, which adds some additional cost, this layer contributes to improved overall performance, as we will show. Please also note that our scheme assumes a constant a fix number of intermediate servers.
In Section 7.1.2, we evaluate our scheme against related works based on essential security and functionality criteria, including threat model support, resistance to model inconsistency attacks, and tolerance for user dropouts.
7.1.1 Computation and Communication Complexity Analysis
In our scheme, each user experiences a computation complexity of $\mathcal{O}(|v|)$, primarily due to the generation of $|v|$ masking parameters for the input vector as shown in Table I. Both the intermediate server and the aggregator have a computation complexity of $\mathcal{O}(m+|v|)$, stemming from the partial and final aggregation of masking parameters from $m$ participating users. As highlighted in Table I, our scheme significantly reduces computation time at both the user and aggregator levels, while introducing a manageable additional $\mathcal{O}(m+|v|)$ complexity for the intermediate server. Notably, this extra complexity is well-suited to the intermediate servers, which are typically much more computationally powerful than users, further enhancing the overall efficiency of our approach.
In our scheme, each user transmits its masked model parameters to the intermediate servers (with the number of intermediaries kept constant) and the aggregator, resulting in a communication complexity of $\mathcal{O}(|v|)$ for the user. Correspondingly, the communication complexity at both the intermediate server and the aggregator is $\mathcal{O}(m\cdot|v|)$. As demonstrated in Table I, our scheme significantly reduces the communication overhead for both the user and aggregator compared to related works, largely by avoiding the use of Shamir’s secret sharing technique and eliminating direct user-aggregator communication during the masking phase. Although our scheme introduces additional communication costs for the intermediate servers, these are manageable given that intermediate servers mainly communicate with the aggregator and are typically more robust in handling higher loads. This design ensures a more scalable and efficient communication process, particularly in environments with large numbers of users. Additionally, our scheme introduces a communication complexity of $\mathcal{O}(n)$ per user due to the need to receive the tuple $\big{<}R,S\big{>}$. This complexity supports our scheme’s unique capability to detect model inconsistency attacks, providing a valuable enhancement in security that is not available in other approaches.
7.1.2 Security and Functionality Comparison
Our scheme offers robust security against both semi-honest and malicious adversaries by carefully configuring system parameters, and it maintains tolerance for user dropouts of up to $m-2$ users at any time. Unlike many existing schemes such as [12], its variants and [19], our approach does not require fixed participation from the start; users can join or dropout at any stage of the training process, enhancing practical applicability. Additionally, our scheme is efficient, requiring only two rounds of communication per iteration and a single setup phase, in contrast to other methods that involve a setup phase in each iteration, leading to increased overhead. Furthermore, our scheme can detect model inconsistency attacks, adding valuable security with a manageable increase in computation and communication costs.
7.2 Experimental Result

### v3

@@ 6.2. Main Protocol
The intermediate servers and aggregator collect masked models from at least $t$ users, where $t$ is the threshold required for each training round. This threshold $t$ is critical for maintaining the security of the masked models. For example, if only two users participate in a round, one user’s local model could be exposed to the other, increasing the risk of collusion. Therefore, the larger the value of $t$, the more secure the system becomes. The intermediate servers and the aggregator wait for a certain timeout period to gather enough masked models before aborting the process if an insufficient number is collected. Users are free to drop out and join at any time.
@@
Afterward, the aggregated global model $\theta$ is shared with all active participants for the next round of training. Simultaneously, the aggregator sends the tuple $V$ to each intermediate server. The intermediate server forwards a part of the tuple $V$ to the users in $\mathrm{I}$ for verification. If the verification of the tuples is successful, the users participate in the next round; otherwise, they abort.
Figure 3 shows two versions of the protocol: one for the semi-honest model and one for the malicious model. In the semi-honest model, where all entities adhere to the protocol without deviation, signatures and the PKI are not needed; these parts are marked in red to indicate they can be skipped. In the malicious model, signatures and the PKI are required to ensure proper authentication, message integrity, and non-repudiation.
7. Security Analysis
In this section, we present the security claims and their respective proofs for our proposed protocol discussed in Section 6.2.
7.1. Semi-Honest Model
In this section, we demonstrate that even if a threshold number of users, denoted as $t_{c}$, and intermediate servers, denoted as $t_{f}$, collude among themselves or with the aggregator, they cannot learn about the remaining genuine users’ local weighted models. We follow a similar security model to those in (et al., 2017) and (Liu et al., 2023).
We consider three scenarios: first, a subset of users is dishonest and collude with adversaries or among themselves, while the intermediate servers and the aggregator remain honest. Second, a subset of intermediate servers is dishonest and can collude, while the users and the aggregator are honest. Third, a subset of users and intermediate servers, along with the aggregator, are dishonest and can collude. Our goal is to show that if fewer than $t_{c}$ (where $t_{c}>2$) and $t_{f}$ (where $t_{f}>1$) entities are compromised, our protocol still protects the individual locally trained models of the remaining genuine users. We assume that $\mathbb{U}$ and $\mathbb{F}$ are the total number of users and intermediate servers, respectively. We also assume that $\mathcal{U}_{i}$ are the number of participating users in the $i^{th}$ round, and $x_{\mathbb{U}}$ is the locally trained models of the users $\mathbb{U}$.
@@ 7.1. Semi-Honest Model
Proof.
We present the detailed security proof in the Appendix C. ∎
Theorem 7.2.
@@
Proof.
We present the detailed security proof in the Appendix D. ∎
Theorem 7.3.
@@
Proof.
We present the detailed security proof in the Appendix E. ∎
7.2. Malicious Model
@@
In this section, we follow the standard security model as in (et al., 2017) and (Liu et al., 2023) to demonstrate that our protocol is secure against active malicious attackers.
@@
In this security model, we need to consider three cases. The first case is the Sybil Attack, where the adversaries can simulate a specific honest user $u_{i}$ to get its inputs. Please note that our protocol is resistant to this attack, as it uses digital signatures to verify the origin of the messages. The second case involves the aggregator sending different lists of participating users in a round to the honest intermediate servers. This could lead to the disclosure of the local weighted models of the honest users. Our protocol detects this type of attack during the model parameter verification phase, where each user can verify the list of participating users in a round received from the intermediate servers. Any difference in the lists will indicate an attack.
@@
The third case involves an aggregator dishonestly dropping honest users at any round. We use the Random Oracle model to prove that our protocol is secure against any malicious dropout of honest users by adversaries. Let’s consider $M_{c}$ as a probabilistic polynomial-time algorithm representing the next message function of participants in $C$, which enables users in $C$ to dynamically select their inputs at any round of the protocol and the list of participating users. The Random Oracle can output the sum of a dynamically selected subset of honest clients for the simulator $\mathrm{Sim}$, which is indistinguishable from the combined view of adversaries in the real protocol execution $\mathsf{Real}(M_{c})$. There are three possible scenarios for this type of active attack: first, the malicious users can collude among themselves; second, the malicious intermediate servers can collude among themselves; and finally, the malicious users and intermediate servers can collude with the aggregator. We present the following three theorems to demonstrate our security proofs.
@@
Table 3. Running Time (ms) for User in Different Phases with a Vector Size of $48k$ and $64$-bit Length.
Number of Users $100$ $200$ $300$ $400$ $500$
Scheme e-SeaFL Ours e-SeaFL Ours e-SeaFL Ours e-SeaFL Ours e-SeaFL Ours
Masking Time $201432$ $2193$ $202927$ $2324$ $210059$ $2487$ $207008$ $2329$ $190656$ $2199$
Masking Time# 202355 $4896$ 209494 $5113$ 212300 $5241$ 205440 $5004$ 207617 $5218$
Model Verification $93657498$ $3864$ 92982871 $3831$ 92806374 $3822$ 91344514 $3845$ 91844664 $3808$
Model Verification# 96533813 $3734$ 96553259 $3767$ 93273166 $3826$ 96467621 $3962$ 95497198 $3958$
Total User Running Time 194063757 $6057$ 185477547 $6155$ 190296465 $6309$ 188311203 $6174$ 188888499 $6007$
Total User Running Time# 195167142 $8630$ 196882805 $8880$ 192445184 $9067$ 194149205 $8966$ 195775738 $9176$
* represent cost in the semi-honest model; # represent cost in the malicious model.
(a) Intermediate Server
(b) Aggregator
@@
Figure 4. Average Running Time of Intermediate Server and Aggregator in the Semi-Honest Model per Round
(a) Intermediate Server
(b) Aggregator
@@
Figure 5. Average Running Time of Intermediate Server and Aggregator in the Malicious Model per Round
(a) Intermediate Server
(b) Aggregator
@@
Figure 6. Average Running Time of Intermediate Server and Aggregator in the Semi-Honest Model per Round
(a) Intermediate Server
(b) Aggregator
@@
Figure 7. Average Running Time of Intermediate Server and Aggregator in the Malicious Model per Round
Theorem 7.4.
@@
Proof.
This proof is similar to that of Theorem 7.1, as the users in $C_{c}$ gain no knowledge about $x_{\mathbb{U}\setminus C_{c}}$. Consequently, the simulator $\mathsf{Sim}$ can assign real inputs to the dishonest users while assigning dummy inputs to the remaining users, accurately replicating the perspectives of the users in $C_{c}$. Thus, the joint views of the users in $C_{c}$ in the simulation are indistinguishable from those in $\mathsf{Real}$. ∎
Theorem 7.5.
@@
Proof.
We present the detailed security proof in the Appendix F. ∎
Theorem 7.6.
@@
Proof.
We present the detailed security proof in the Appendix G. ∎
8. Experimental Evaluation
In this section, we present experimental results to demonstrate the practical efficiency of our proposed protocol. A theoretical comparison of functionality and security is provided in Table 1, while the complexity analysis of communication and computation costs is shown in Table 2. These aspects were briefly discussed in Section 2. Overall, the theoretical findings indicate that our protocol offers enhanced functionality and stronger security guarantees, while maintaining communication and computation costs comparable to e-SeaFL, which is the most efficient single-setup-based secure aggregation protocol to date.
To further assess practicality, we conduct experiments comparing our protocol with e-SeaFL, focusing on concrete running times for the user, intermediate server, and aggregator. This evaluation highlights the computational efficiency of our protocol. Additionally, we report the accuracy achieved on the MNIST (Lecun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets, comparing our results with those obtained using the standard FedAvg method (McMahan et al., 2017) (a baseline FL approach).